Prelab Activity:

Titrations Continued – Titration of Household Products

  1. A 2.40 g sample of vinegar was added to an Erlenmeyer flask along with 100 mL of deionized water and 3 drops of phenolphthalein indicator. It took 22.15 mL of 0.0981 M NaOH (aq) to reach the faint pink endpoint. The following balanced chemical equation represents the chemistry of the titration:

NaOH (aq) + CH3COOH (aq) à CH3COONa+ (aq) + H2O (l)

  1. Calculate the mass % of acetic acid (CH3COOH) in the vinegar sample.
  2. Two subsequent trials found the mass % of acetic acid in vinegar to be 5.54% and 5.53%, respectively. Using the mass % found in part a, along with the values for subsequent trials, report an average value and standard deviation of mass % for the three trials. Report answer in the format: AVG ± S.D.

2.4.0 mL of 3.00 M HCl reacts completely with the CaCO3 (active ingredient) in an antacid tablet. The remaining solution (excess HCl) is then titrated with 0.0981 M NaOH to the bromothymol blue endpoint.

  1. Write the balanced chemical equation for the reaction of CaCO3 (s) with HCl (aq).
  1. If the volume of NaOH delivered at the endpoint is 22.20 mL, determine the amount (in milligrams) of CaCO3 that was contained in the antacid tablet.
Answer #1

1-

The given neutralization reaction is-

NaOH (aq) + CH3COOH (aq) ---------> CH3COO–Na+ (aq) + H2O (l)

That means 1 mole of NaOH is nedded to neutralize 1 mole of CH3COOH

Now given

volume of NaOH used for the reaction = 22.15 mL

Concentration of NaOH used for the reaction = 0.0981 M

That means moles of NaOH used for the reaction = concentration * volume

= 0.0981 M * 22.15 mL

= 0.0981 mole / 1000 mL * 22.15 mL

= 0.00217 moles

That means moles of CH3COOH neutralized = 0.00217 moles

So we can say in the taken vinegar sample, we have 0.00217 moles of CH3COOH

So mass of CH3COOH present = moles * molar mass of CH3COOH

= 0.00217 moles * 60 g/mole

= 0.13 g

So mass % of acetic acid (CH3COOH) in the vinegar sample

= (mass of acetic acid present / mass of vinegar sample) * 100

= (0.13 g / 2.40 g) * 100

= 5.43 %

b-

Now given in subsequent 2 trials, % of acetic acid (CH3COOH) found = 5.54% and 5.53%

So average % of acetic acid (CH3COOH) found = (5.54% + 5.53% + 5.43 % ) / 3

= 5.5 %

Now we can calcuclate standard deviation as-

i- Caluclate the mean = 5.5

ii- substract the mean from each value i.e

5.54 - 5.5 = 0.04

5.53 - 5.5 = 0.03

5.43 - 5.5 = -0.07

iii- square each of the deviation values and add them i.e

(0.04)2 +  (0.03)2 +  (-0.07)2 = 0.0074‬

iv- substract 1 from total number of data sets. We have total number of data sets = 3 (i.e 5.54, 5.43)

So here 3-1 = 2

v- devide the sum of squares by this value found in step iv i,e

0.0074‬ / 2 = 0.0037‬

vi- Tkae the square root of the numebr i.e

sq. rt (0.0037) = 0.061

This is the standard deviation.

So the mass % of acetic acid (CH3COOH) found = Avg ± SD

= 5.5 % ± 0.061